OpenStudy (anonymous):

Can the set of lengths be the side lengths of a right triangle? 8ft 15ft 17ft A. no B. yes

OpenStudy (anonymous):

Try them out. use pythagorean theorem.

OpenStudy (anonymous):

I did and it makes no since

OpenStudy (campbell_st):

welll you are being asked if \[8^2 + 15^2 = 17^2\] or does \[64 + 225 = 289\] that's all you need to do... then make a decision as to whether the number sentence is correct

OpenStudy (anonymous):

so its yes?